# Relation between equilibrium quantum phase transitions and dynamical quantum phase transitions in two-band systems

### v1

@@ III The model and results
Figure 6: (a) The image of ¯$\bar{\mathcal{L}}_{k}$ versus $\mathcal{F}_{k}^{\text{q}}$. (b)-(e) are images of $\mathcal{F}_{k}^{\text{q}}$ and ¯$\bar{\mathcal{L}}_{k}$ versus $k$. The dark blue solid lines are $\mathcal{F}_{k}^{\text{q}}$. The light blue solid lines are ¯$\bar{\mathcal{L}}_{k}$. The blue horizontal dashed lines represent 2/$\sqrt{2}/2$. The black vertical dashed lines denote corresponding $k_{c}$. Post-quench parameters are: (b) $h_{f}=-2,\ \eta_{f}=-2$; (c) $h_{f}=-1.1,\ \eta_{f}=-2$; (d) $h_{f}=0,\ \eta_{f}=-2$; (e) $h_{f}=2,\ \eta_{f}=-2$. Here $h_{i}=-2$ and $\eta_{i}=0.8$.
@@ IV Conclusion and discussion
In summary, we proposed the concept of the quench fidelity and applied it to unveil the relation between EQPTs and DQPTs from a new perspective. By using ¯$\bar{\mathcal{L}}_{k}$ to denote the minimum of each $\mathcal{L}_{k}(t)$ in time, and defining the quench fidelity $\mathcal{F}^{\text{q}}$ as the module of the overlap of ground states of the Hamiltonian with pre-quench parameter and the one with post-quench parameter, we found the relation between ¯$\bar{\mathcal{L}}_{k}$ and $\mathcal{F}_{k}^{\text{q}}$, which reveals that the occurrence of DQPTs corresponds to $k_{c}$-modes of $\mathcal{F}^{\text{q}}$ which satisfy $\mathcal{F}_{k_{c}}^{\text{q}}=\sqrt{2}/2$. Therefore, a necessary and sufficient condition for the existence of DQPTs is the quench fidelity of the pre-quench system and the post-quench system has at least a $k_{c}$-mode which satisfies $\mathcal{F}_{k_{c}}^{\text{q}}=\sqrt{2}/2$.
@@
To demonstrate our theory, we took the one-dimensional spin-$\frac{1}{2}$ anisotropic XY model as a concrete example and calculated it in detail. According to the number $n_{k_{c}}$ of $k_{c}$-modes, we divided the post-quench parameter space into three kinds of regions, and thus got the dynamical phase diagram of DQPTs of the XY model for given pre-quench parameters. In comparison with the original equilibrium phase diagram, the dynamical phase diagram of XY model has two additional boundaries, which demonstrates the essential difference between EQPTs and DQPTs. According to the number $n_{k_{0}}$ of $k_{0}$-modes and the number $n_{k_{1}}$ of $k_{1}$-modes, we divided the post-quench parameter space into different regions. Although phase diagrams of $k_{0}$-modes and $k_{1}$-modes are a little different from the original equilibrium phase diagram, the overlapping boundaries of them match exactly the equilibrium phase boundaries, indicating that there are changes in the values of both $n_{k_{0}}$ and $n_{k_{1}}$ at equilibrium phase boundaries. Besides, we found that the phase diagram of $k_{0}$-modes has two extra boundaries across which there will appear more or less $k_{0}$-modes, whereas abrupt change in each $\mathcal{F}_{k}^{\text{q}}$ only occurs at equilibrium phase boundaries. The discrepancy between the extra boundaries and equilibrium phase boundaries can be demonstrated by the image of the derivative of the decay rate function of the quench fidelity. Through the image of ¯$\bar{\mathcal{L}}_{k}$ versus $\mathcal{F}_{k}^{\text{q}}$, we concluded that a sufficient condition for the occurrence of DQPTs is the quench fidelity has at least a $k_{0}$-mode and a $k_{1}$-mode simultaneously, whereas DQPTs may or may not occur if the quench fidelity has only $k_{1}$-modes or only $k_{0}$-modes. It means that quenching across the equilibrium phase point is not sufficient for the occurrence of DQPTs unless the quench fidelity owns at least a $k_{1}$-mode, and quenching within the same equilibrium phase can also give rise to DQPTs, which accounts for the mismatch between dynamical phases and equilibrium phases.
@@
Our work clarifies clearly the relation between EQPTs and DQPTs and provides a new insight for future study on conditions for the occurrence of DQPTs. Nonetheless, our theory is currently limited to two-band systems, and further researches on promoting it to a broader range of models are anticipated.
@@ Appendix A The detailed derivation of formulas
Assume the argument of $\left\langle 1_{i}|0_{f}\right\rangle\left\langle 0_{f}|0_{i}\right\rangle$ is $\theta$, then $\left\langle 1_{i}|0_{f}\right\rangle\left\langle 0_{f}|0_{i}\right\rangle=e^{i\theta}\left|\left\langle 1_{i}|0_{f}\right\rangle\left\langle 0_{f}|0_{i}\right\rangle\right|=e^{i\theta}/2$, thus we can derive

### v2

@@ III The model and results
Figure 6: (a) The image of $\bar{\mathcal{L}}_{k}$ versus $\mathcal{F}_{k}^{\text{q}}$. (b)–(e) are images of $\mathcal{F}_{k}^{\text{q}}$ and $\bar{\mathcal{L}}_{k}$ versus $k$. The dark blue solid lines are $\mathcal{F}_{k}^{\text{q}}$. The light blue solid lines are $\bar{\mathcal{L}}_{k}$. The blue horizontal dashed lines represent $\sqrt{2}/2$. The black vertical dashed lines denote corresponding $k_{c}$. Postquench parameters are (b) $h_{f}=-2,\ \eta_{f}=-2$; (c) $h_{f}=-1.1,\ \eta_{f}=-2$; (d) $h_{f}=0,\ \eta_{f}=-2$; (e) $h_{f}=2,\ \eta_{f}=-2$. Here $h_{i}=-2$ and $\eta_{i}=0.8$.
@@ IV Conclusion and discussion
In summary, we proposed the concept of the quench fidelity and applied it to unveil the relation between EQPTs and DQPTs from a new perspective. By using $\bar{\mathcal{L}}_{k}$ to denote the minimum of each $\mathcal{L}_{k}(t)$ in time, and defining the quench fidelity $\mathcal{F}^{\text{q}}$ as the module of the overlap of ground states of the Hamiltonian with the prequench parameter and the one with the postquench parameter, we found the relation between $\bar{\mathcal{L}}_{k}$ and $\mathcal{F}_{k}^{\text{q}}$, which reveals that the occurrence of DQPTs corresponds to $k_{c}$-modes of $\mathcal{F}^{\text{q}}$ which satisfy $\mathcal{F}_{k_{c}}^{\text{q}}=\sqrt{2}/2$. Therefore, a necessary and sufficient condition for the existence of DQPTs is that the quench fidelity of the prequench system and the postquench system has at least a $k_{c}$-mode which satisfies $\mathcal{F}_{k_{c}}^{\text{q}}=\sqrt{2}/2$.
@@
To demonstrate our theory, we took the one-dimensional spin-$\frac{1}{2}$ anisotropic XY model as a concrete example and calculated it in detail. According to the number $n_{k_{c}}$ of $k_{c}$-modes, we divided the postquench parameter space into three kinds of regions, and thus got the dynamical phase diagram of DQPTs of the XY model for given prequench parameters. In comparison with the original equilibrium phase diagram, the dynamical phase diagram of the XY model has two additional boundaries, which demonstrates the essential difference between EQPTs and DQPTs. According to the number $n_{k_{0}}$ of $k_{0}$-modes and the number $n_{k_{1}}$ of $k_{1}$-modes, we divided the postquench parameter space into different regions. Although phase diagrams of $k_{0}$-modes and $k_{1}$-modes are a little different from the original equilibrium phase diagram, the overlapping boundaries of them match exactly the equilibrium phase boundaries, indicating that there are changes in the values of both $n_{k_{0}}$ and $n_{k_{1}}$ at equilibrium phase boundaries. Besides, we found that the phase diagram of $k_{0}$-modes has two extra boundaries across which there will appear more or fewer $k_{0}$-modes, whereas abrupt change in each $\mathcal{F}_{k}^{\text{q}}$ only occurs at equilibrium phase boundaries. The discrepancy between the extra boundaries and equilibrium phase boundaries can be demonstrated by the image of the derivative of the decay rate function of the quench fidelity. Through the image of $\bar{\mathcal{L}}_{k}$ versus $\mathcal{F}_{k}^{\text{q}}$, we concluded that a sufficient condition for the occurrence of DQPTs is that the quench fidelity has at least a $k_{0}$-mode and a $k_{1}$-mode simultaneously, whereas DQPTs may or may not occur if the quench fidelity has only $k_{1}$-modes or only $k_{0}$-modes. This means that quenching across the equilibrium phase point is not sufficient for the occurrence of DQPTs unless the quench fidelity owns at least a $k_{1}$-mode, and quenching within the same equilibrium phase can also give rise to DQPTs, which accounts for the mismatch between dynamical phases and equilibrium phases.
@@
Our work clarifies clearly the relation between EQPTs and DQPTs and provides a new insight for future study on conditions for the occurrence of DQPTs. Nonetheless, our theory is currently limited to two-band systems, and further research on promoting it to a broader range of models is anticipated.
@@ Appendix A The detailed derivation of formulas
Without loss of generality, we assume that the argument of $\left\langle 1_{i}|0_{f}\right\rangle\left\langle 0_{f}|0_{i}\right\rangle$ is $\theta$, then $\left\langle 1_{i}|0_{f}\right\rangle\left\langle 0_{f}|0_{i}\right\rangle=e^{i\theta}\left|\left\langle 1_{i}|0_{f}\right\rangle\left\langle 0_{f}|0_{i}\right\rangle\right|=e^{i\theta}/2$, and thus we can derive